# Beyond Memorization: A Rigorous Evaluation Framework for Medical Knowledge Editing

### v2

@@ 4.3 Evaluation Metrics
As shown in Figure 2, the knowledge editing is conducted with questions $q\in\mathcal{Q}_{\mathrm{ori}}$ and their corresponding knowledge $k$ defined in each editing paradigm via Equation 1 to inject knowledge into LLM. The edited model $\theta^{\prime}$ is then evaluated on three test sets, i.e., $q\in\mathcal{Q}_{\mathrm{ori}}\cup\mathcal{Q}_{\mathrm{gen}}\cup\mathcal{Q}_{\mathrm{ret}}$ to measure the efficacy, generalization, and retention of each editing method with the accuracy of the predicted answer offered by the edited LLM.
To further evaluate the interpretability of the edited model, we also instruct the model to provide a rationale for its final answer, which is compared with the ground-truth interpretation using ROUGE-L (Lin, 2004) and BLEU scores (Papineni et al., 2002). Detailed calculations are provided in Section C.4.
@@ 4.4 Editing Method Selection
We choose six representative editing methods from three categories: Fine-Tuning methods - LoRA (Hu et al., 2022), Parameter-Modifying methods - ROME (Meng et al., 2022) and MEMIT (Meng et al., ), and Parameter-Preserving methods - GRACE (Hartvigsen et al., 2023). Moreover, we include two newer Parameter-Modifying variants: AnyEdit (Jiang et al., 2025) and AlphaEdit (Fang et al., 2024), which are state-of-the-art approaches supporting long-form and continual knowledge editing, respectively. We exclude retrieval-based editing methods (Song et al., 2024; Shi et al., 2024; Chen et al., 2024b; Zheng et al., 2023) since they do not directly update knowledge within the model parameters. Details of the editing methods are provided in Appendix B.
Method Metric Pre-Edit LLaMA-8B LLaMA-3B
MedExQAedit MedMCQAedit MedExQAedit MedMCQAedit
LoRA Eff. 0 43.5 46.6 36.7 14.4
Gen. 0 41.3 41.6 43.3 33.7
Ret. 100 63.0 70.8 63.3 52.9
[…]
ROME Eff. 0 37.0 32.7 23.3 25.7
Gen. 0 43.5 29.6 26.7 25.1
Ret. 100 63.0 61.6 56.7 56.1
[…]
MEMIT Eff. 0 39.1 28.3 28.3 16.0
Gen. 0 50.0 25.2 31.7 25.1
Ret. 100 50.0 64.8 63.3 54.5
[…]
GRACE Eff. 0 34.8 36.0 33.3 26.7
Gen. 0 21.7 29.2 10.0 7.0
Ret. 100 76.1 80.7 88.3 93.0
[…]
AnyEdit Eff. 0 34.8 36.6 35.0 23.0
Gen. 0 28.3 25.5 28.3 23.0
Ret. 100 78.3 78.9 71.7 78.1
[…]
AlphaEdit Eff. 0 47.8 43.9 35.0 32.6
Gen. 0 32.6 31.2 31.7 27.8
Ret. 100 76.1 86.7 73.3 77.5
[…]
@@
4.5 Experimental Setup
We evaluate all editing methods on two LLMs: LLaMA-3.1-8B-Instruct and LLaMA-3.2-3B-Instruct [^2] https://llama.meta.com/lama3/. Editing pipelines are implemented based on the EasyEdit framework (Wang et al., 2023). Detailed experimental settings are provided in Appendix C. To ensure interpretable reasoning, all post-edit predictions follow a two-step format: a rationale followed by the final answer. This approach encourages the model to conduct deep reasoning, resulting in both improved prediction accuracy and enhanced interpretability. The output format is illustrated in Section D.4.
5 Main Experiments
@@
  • RQ1: How do current model editing methods perform in the medical domain?
  • RQ2: How do different editing paradigms impact the effectiveness of medical knowledge editing?
  • RQ3: How does the medical knowledge stored in LLMs?
  • RQ4: How does the knowledge in LLMs evolve with sequential editing?
@@ 5.1 Evaluating Editing Methods in Medical Domain (RQ1)
Existing editing methods follow the ground-truth answer editing (GTA-Edit) paradigm, where the model is updated by injecting the answer $a$ as the target knowledge (Meng et al., 2022, ). We follow this paradigm to update LLMs with the ground-truth answer $a$ for each question $q\in\mathcal{Q}_{\mathrm{ori}}$. Then, the model’s performance is evaluated on the three test sets $\mathcal{Q}_{\mathrm{ori}}\cup\mathcal{Q}_{\mathrm{gen}}\cup\mathcal{Q}_{\mathrm{ret}}$, and the question answering accuracy is reported in Table 1. From the results, we observe that:
@@
The efficacy of nearly all methods is below 50%, except for AlphaEdit and LoRA on LLaMA-8B, which achieve only 53.9% and 53% on MedMCQAedit, respectively. This sharply contrasts with previous reports of over 90% efficacy in general domain benchmarks (Meng et al., 2022; Fang et al., 2024), highlighting a significant gap when applying knowledge editing methods to the intricate medical domain.
@@
This finding is reflected in the low generalization scores and noticeable drops in retention. For example, on MedMCQAedit, LoRA achieves only 41.6% (8B) and 33.7% (3B) generalization, while retaining 70.8% and 52.9% of previously correct answers. Despite its strong retention, GRACE shows limited generalization ability, making it less suitable for adapting to new clinical contexts. These results suggest that existing editing paradigm (i.e., GTA-Edit) often leads to surface-level memorization rather than meaningful internalization of medical knowledge. Detailed case analyses are can be found in Appendix E.
@@ 5.2 Evaluation of Editing Paradigms (RQ2)
Figure 4: Interpretability comparison for SGR-Edit and RE-Edit.
This section investigates how different editing paradigms (i.e., GTA-Edit, RE-Edit, and SGR-Edit), affect the performance of existing editing methods on MedMCQAedit. Our key findings are as follows:
RQ2-F1: SGR-Edit yields the highest editing performance.
As shown in Figure 3, SGR-Edit achieves the best performance for the five selected editing methods in both LLaMA-8B and LLaMA-3B, demonstrating that the reasoning knowledge drives is more effective for medical knowledge editing. On LLaMA-8B, AlphaEdit’s average score increases from 53.9% (GTA-Edit) to 55.9% (RE-Edit, +2.0%) and 58.0% (SGR-Edit, +4.1%). LoRA exhibits a similar enhancement, rising from 53.0% (GTA-Edit) to 57.8% (RE-Edit, +4.8%) and 62.5% (SGR-Edit, +9.5%). Overall, SGR-Edit boosts performance by an average of 8.6% relative to GTA-Edit on LLaMA-8B, and by 12.6% on LLaMA-3B, demonstrating its consistently stronger efficacy across model sizes. Importantly, SGR-Edit achieves these gains with only one additional reasoning-generation step, making it both highly effective and readily deployable in practical settings. Full results and further analysis are provided in Appendix D.1.
RQ2-F2: SGR-Edit enables better reasoning interpretation.
We report the rational interpretability of SGR-Edit and RE-Edit in Figure 4. The results indicate that SGR-Edit provides a more comprehensive understanding of the underlying medical knowledge, enabling better reasoning interpretation than RE-Edit. Detailed results and analyses are provided in Appendix D.2.
5.3 Localization of Medical Knowledge in LLMs (RQ3)
Figure 5: Layer-Wise editing performance
This section investigates the layer-wise storage of medical knowledge in LLMs. We edit knowledge across four disjoint layers to check the storage of certain knowledge: For LLaMA-8B, we select layers 4–8, 11–15, 18–22, and 25–29; for the smaller LLaMA-3B, we use layers 4–8, 10–14, 16–20, and 22–26. We sample 100 QA pairs from MedMCQAedit and group target knowledge by token length to assess the impact of knowledge granularity. GTA targets are the shortest answers (<10 tokens), RE explanations span 50–100 or 100–150 tokens, and SGR rationales range from 150–200 to 200–250 tokens. Based on Figure 5, we draw the following findings:
RQ3-F1: Medical knowledge is primarily stored in shallower layers.
For both LLaMA-8B and 3B, editing operations targeting shallower layers 4-8 consistently yield the highest effectiveness. When editing on LLaMA-8B, average scores in layers 4-8 range from 0.433 to 0.583, and for LLaMA-3B, they span from 0.333 to 0.533, consistently outperforming edits in deeper layers.
RQ3-F2: Across all editing paradigms, the most effective edits are consistently achieved in the shallower layers.
Model #Edits Overall $\Delta$Overall Health $\Delta$Health Non-Health $\Delta$Non-Health
LLaMA-8B 0 (Raw) 67.9 70.7 68.0
50 66.4 –1.5 68.5 –2.2 66.6 –1.5
100 63.2 –4.6 61.1 –9.6 64.7 –3.3
LLaMA-3B 0 (Raw) 60.7 63.7 60.5
50 59.1 –1.6 60.7 –3.0 59.1 –1.3
100 56.8 –3.9 54.4 –9.3 57.4 –3.1
Table 2: Sequential editing performance and absolute drops $\Delta$ on MMLU (Accuracy %).
5.4 Reliability of Sequential Medical Edits (RQ4)
Figure 6: Performance over sequential medical editing.
In this section, we study the reliability of sequential edits to medical knowledge in LLMs, where the knowledge is sequentially updated through multiple edits. We assess three representative sequential editing methods: AlphaEdit, LoRA, and MEMIT, through up to 100 sequential edits using samples from MedMCQAedit. We track average editing success rates at edit steps {1, 5, 10, 20, 30, 50, 100}. To better quantify the side effects of medical knowledge editing on the model’s retention of other information, we analyze performance drops in both health and non-health domains of the MMLU benchmark (Hendrycks et al., 2020) after 50 and 100 sequential edits using AlphaEdit.
RQ4-F1: AlphaEdit supports stable and effective sequential editing far better than other methods.
Figure 6 shows that on LLaMA-8B, AlphaEdit maintains an average editing score above 60% through 30 sequential edits, declining moderately to 58% at 50 edits and 41% at 100 edits. In contrast, LoRA drops sharply from 63% after the first edit to just 9% at 50 edits and collapses to 2% at 100 edits. MEMIT falls to near zero by the 50th edit. A comparable pattern is observed on LLaMA-3B: AlphaEdit sustains 46% average performance after 50 edits, while LoRA—despite strong single-edit results, it fails to withstand continual editing, and MEMIT degrades rapidly under sequential editing.
RQ4-F2: Sequential medical edits deteriorate the knowledge in other domains.
As shown in Table 2, after 100 sequential edits on LLaMA-8B, health-domain accuracy on MMLU declines more sharply (e.g., $\Delta=-9.6$% on LLaMA-8B, $\Delta=-9.3$% on LLaMA-3B), which is expected given the exclusively medical nature of the edit data. However, non-health accuracy falls from 68.0% to 64.7% ($\Delta=-3.3$%), and on LLaMA-3B from 60.5% to 57.4% ($\Delta=-3.1$%), suggesting that continual updates to medical knowledge can lead to the erosion of out-of-domain knowledge.
@@ 6 Conclusion
In this paper, we introduce MedEditBench, a rigorous evaluation framework for assessing knowledge editing methods in the medical domain. It features diverse metrics and editing paradigms to examine how different knowledge sources impact editing performance. Our experimental results reveal that the current editing paradigm updating with a simple final answer, tend to induce superficial memorization rather than foster genuine understanding of the underlying rationale. To overcome this limitation, we propose the self-generated rationale editing (SGR-Edit), a novel editing paradigm where the model first generates an evidence-grounded rationale as the editing knowledge target. SGR-Edit demonstrates remarkable improvements in medical knowledge editing, allowing deeper internalization of new knowledge and generalize edited LLM to new scenarios. Additionally, our analysis uncovers that medical knowledge is usually stored in shallower LLM layers, and when sequential editing on medical knowledge, the LLM typically suffers from degradation of broader knowledge. This finding provides practical guidance for medical applications and further research in this area.
@@
  • Evaluation Coverage: Due to limited GPU resources, our main experiments focus on six representative editing methods applied to two widely used LLaMA models. To broaden coverage, we additionally report single-edit results on Qwen2.5-7B in Appendix D.3.

### v3

@@ 4.3 Evaluation Metrics
As shown in Figure 2, the knowledge editing is conducted with questions $q\in\mathcal{Q}_{\mathrm{ori}}$ and their corresponding knowledge $k$ defined in each editing paradigm via Equation˜1 to inject knowledge into LLM. The edited model $\theta^{\prime}$ is then evaluated on three test sets, i.e., $q\in\mathcal{Q}_{\mathrm{ori}}\cup\mathcal{Q}_{\mathrm{gen}}\cup\mathcal{Q}_{\mathrm{ret}}$ to measure the efficacy, generalization, and retention of each editing method with the accuracy of the predicted answer offered by the edited LLM.
To further evaluate the interpretability of the edited model, we also instruct the edited LLM to provide a rationale for the final answer, which is compared with the ground-truth interpretation using ROUGE-L (Lin, 2004), BLEU scores (Papineni et al., 2002), and Quality of Reasoning (QoR) scored via LLM-as-a-Judge. Detailed calculations are provided in Section˜C.4.
@@ 4.4 Editing Method Selection
We choose six representative editing methods from three categories: Fine‐Tuning methods - LoRA (Hu et al., 2022), Parameter‐Modifying methods - ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), and Parameter‐Preserving methods - GRACE (Hartvigsen et al., 2023). Moreover, we include two newer Parameter‐Modifying variants: AnyEdit (Jiang et al., 2025) and AlphaEdit (Fang et al., 2025), which are state-of-the-art approaches supporting long-form and continual knowledge editing, respectively. Details of the KE methods are provided in Appendix B.
4.5 Experimental Setup
We evaluate KE methods on three LLMs: LLaMA-3.1-8B-Instruct, LLaMA-3.2-3B-Instruct [^1] https://huggingface.co/meta-llama, and Qwen2.5-7B-Instruct [^2] https://huggingface.co/Qwen. We utilize layers [4, 5, 6, 7, 8] for editing based on our findings from detailed experiments in Appendix D.1. Detailed experimental settings are provided in Appendix C.
5 Main Experiments
Method Metric Pre-Edit LLaMA-8B LLaMA-3B Qwen-7B
MedExQAedit MedMCQAedit MedExQAedit MedMCQAedit MedExQAedit MedMCQAedit
LoRA Eff. 0 43.5 46.6 36.7 14.4 66.0 44.0
Gen. 0 41.3 41.6 43.3 33.7 48.0 36.0
Ret. 100 63.0 70.8 63.3 52.9 84.0 77.0
avg. \ 49.3 53.0 47.8 33.7 66.0 52.3
ROME Eff. 0 37.0 32.7 23.3 25.7 47.8 50.0
Gen. 0 43.5 29.6 26.7 25.1 56.5 19.3
Ret. 100 63.0 61.6 56.7 56.1 58.7 56.0
avg. \ 47.8 41.3 35.6 35.7 54.3 41.8
MEMIT Eff. 0 39.1 28.3 28.3 25.0 45.7 47.0
Gen. 0 50.0 25.2 31.7 23.0 54.3 30.0
Ret. 100 50.0 64.8 63.3 48.5 63.0 57.0
avg. \ 46.4 39.4 41.1 32.2 54.3 44.7
GRACE Eff. 0 34.8 36.0 33.3 26.7 39.1 23.0
Gen. 0 21.7 29.2 10.0 7.0 43.5 32.0
Ret. 100 76.1 80.7 88.3 93.0 63.0 69.0
avg. \ 44.2 48.7 43.9 42.2 48.6 41.3
AnyEdit Eff. 0 34.8 36.6 35.0 23.0 39.1 26.0
Gen. 0 28.3 25.5 28.3 23.0 52.2 27.0
Ret. 100 78.3 78.9 71.7 78.1 73.9 83.0
avg. \ 47.1 47.0 45.0 41.4 55.1 45.3
AlphaEdit Eff. 0 47.8 43.9 35.0 32.6 60.0 51.0
Gen. 0 32.6 31.2 31.7 27.8 56.7 42.0
Ret. 100 76.1 86.7 73.3 77.5 66.7 64.0
avg. \ 52.2 53.9 46.7 46.0 61.1 52.3
@@ 5 Main Experiments
RQ1: How do current KE methods perform in the medical domain?
RQ2: How do different KE paradigms impact the editing performance?
RQ3: How does sequential medical KE affect the reliability and stability of in-domain knowledge?
RQ4: Whether sequential medical KE affects broader domain knowledge?
RQ5: How does sequential medical KE affect general LLM capabilities?
@@ 5.1 Evaluating Editing Methods in Medical Domain (RQ1)
Existing editing methods follow the ground‐truth answer editing (GTA‐Edit) paradigm, where the model is updated by injecting the answer $a$ as the target knowledge (Meng et al., 2022, 2023). We follow this paradigm to update LLMs with the ground‐truth answer $a$ for each question $q\in\mathcal{Q}_{\mathrm{ori}}$. Then, the model’s performance is evaluated on the three test sets $\mathcal{Q}_{\mathrm{ori}}\cup\mathcal{Q}_{\mathrm{gen}}\cup\mathcal{Q}_{\mathrm{ret}}$. From the results in Table˜1, we observe that:
@@
The efficacy of nearly all methods is below 50%. This sharply contrasts with previous reports of over 90% efficacy in general domain benchmarks (Meng et al., 2022; Fang et al., 2025), highlighting a significant gap when applying current KE methods to the intricate medical domain.
@@
This finding is reflected in the low generalization scores and noticeable drops in retention. For example, on MedMCQAedit, LoRA achieves only 41.6% (8B) and 33.7% (3B) generalization, while retaining 70.8% and 52.9% of previously correct answers. Despite its strong retention, GRACE shows limited generalization ability, making it less suitable for adapting to new clinical contexts. These results suggest that existing editing paradigm (i.e., GTA-Edit) often leads to surface-level memorization rather than meaningful internalization of medical knowledge. Detailed case analyses can be found in Appendix E.
@@ 5.2 Evaluation of Editing Paradigms (RQ2)
This section investigates how different editing paradigms (i.e., GTA-Edit, RE-Edit, and SGR-Edit), affect the performance of existing editing methods. To ensure the factual grounding of the rationales generated by the base LLM, we performed a targeted evaluation on a subset to assess rational factuality through a rigorous, two-stage process of AI-judge scoring and human verification (see Appendix D.2). Our key findings are as follows:
RQ2-F1: SGR-Edit yields the highest editing performance.
As shown in Figure 3, SGR-Edit achieves the best performance for the five selected KE methods across various base LLMs. Overall, SGR-Edit boosts performance by an average of 7.7% relative to GTA-Edit on LLaMA-8B, by 11.3% on LLaMA-3B, and by 6.3% on Qwen-7B, demonstrating its consistently stronger capability across different model sizes and architectures. Importantly, SGR-Edit achieves these gains with only one additional reasoning‐generation step, making it both highly effective and readily deployable. We discussed why SGR-Edit is practical in real-world employment in Appendix D.3. We also experimented with externally generated rationales in Appendix D.4. The complete results and further analysis of various editing paradigms are provided in Appendix D.5.
@@
Figure 4: Interpretability comparison for SGR-Edit and RE-Edit. Quality of Reasoning (QoR) is a normalized (0–1) score via an AI judge (DeepSeek-V3 (DeepSeek-AI, 2024)), assessing the generated rationales across five dimensions: factual accuracy, logical flow, relevance, completeness, and answer correctness (see prompt in Figure˜14).
RQ2-F2: SGR-Edit enables better reasoning interpretability.
We compare the knowledge quality enabled by SGR-Edit and RE-Edit using MEMIT in Figure 4. SGR-Edit consistently and substantially outperforms RE-Edit across all models and metrics, confirming that using the rationale-derived knowledge for editing results in deeper knowledge internalization. Full results for three KE methods are detailed in Appendix D.6.
5.3 Sequential Editing Impacts on In-Domain Knowledge (RQ3)
We investigate how sequential editing affects the in-domain knowledge in LLMs, where knowledge items are continuously updated through multiple editing steps. Performance is tracked over {1, 5, 10, 20, 30, 50, 100} sequential edits using samples from MedEditBench, as shown in Figure 5. We report our findings from three perspectives: model-wise, method-wise, and paradigm-wise:
RQ3-F1: Qwen experiences the most severe degradation under sequential edits, while LLaMA models exhibit greater stability.
All models experience gradual declines as the number of edits increases. However, the performance drop on Qwen-7B is substantially larger than that on LLaMA-8B or LLaMA-3B.
RQ3-F2: Despite strong single-edit performance, LoRA suffers the sharpest decline under sequential edits, whereas AlphaEdit remains the most stable.
Across all models and paradigms (best viewed in color), LoRA quickly collapses after several edits, showing near-zero accuracy after 50–100 updates. MEMIT also degrades rapidly, while AlphaEdit sustains moderate performance even after 100 edits, highlighting its robustness in sequential editing.
RQ3-F3: Among editing paradigms, SGR-Edit is the most resistant to sequential degradation.
The rationale-driven SGR-Edit paradigm consistently preserves higher accuracy over repeated updates compared with GTA-Edit and RE-Edit (best viewed in line and marker). This indicates that SGR-Edit enables deeper knowledge integration and reduces catastrophic forgetting during continuous medical knowledge updates.
Figure 5: Sequential editing performance on the internal benchmark.
5.4 Sequential Editing Impacts on External Domain Knowledge (RQ4)
After assessing internal medical edits, we further investigate their influence on external-domain knowledge, covering both medical and non-medical aspects. We split the MMLU benchmark accordingly (see Appendix D.8 for subject mapping) to examine whether sequential medical editing propagates effects beyond the edited domain. Observations from Figure 6 are:
Figure 6: Sequential editing performance on the external MMLU benchmark.
RQ4-F1: Sequential medical edits consistently degrade both medical and non-medical knowledge from the external domain.
After 100 sequential edits using AlphaEdit on the internal MedEditBench, accuracy drops substantially not only in the medical domain from the MMLU benchmark — especially for Qwen-7B (-28.3%) compared to LLaMA-8B (-9.5%) and LLaMA-3B (-9.3%), but also in non-medical categories (-3.3% on LLaMA-8B and -3.1% on LLaMA-3B). This indicates that sequential editing in internal medical knowledge erodes broader knowledge.
RQ4-F2: Advanced editing method mitigates external knowledge deterioration.
While both KE methods degrade performance, AlphaEdit consistently retains higher general-domain accuracies, as indicated by the solid lines, demonstrating that sequential optimization from the AlphaEdit mitigates catastrophic forgetting across domains. We report category-wise results on MMLU general-domain tasks in Appendix Table 8. We also provide detailed sequential editing results for traditional methods like ROME in Table 10, which exhibits similar catastrophic forgetting patterns to MEMIT, further confirming the necessity of sequential-optimized methods. Further detailed comparison of different editing paradigms (GTA-Edit, RE-Edit, and SGR-Edit) and SGR edits with varying target knowledge lengths is presented in Table 9. We find that the primary driver of broader domain knowledge deterioration is the KE method (e.g., MEMIT’s severe drop vs. AlphaEdit’s minor drop), rather than the specific editing paradigm. Crucially, increasing the target knowledge length in SGR-Edit does not cause significantly greater external knowledge degradation.
5.5 Sequential Editing Impacts on General LLM Capabilities (RQ5)
Table 2: Comparison of sequential editing performance on general benchmarks (AlphaEdit vs. MEMIT).
Model Benchmark # Edits (AlphaEdit) # Edits (MEMIT)
0 50 100 0 50 100
LLaMA-8B HellaSwag 79.2 78.8 78.5 79.2 26.2 25.9
TriviaQA 51.9 50.1 48.4 51.9 0.0 0.0
TruthfulQA 54.1 55.1 54.9 54.1 47.6 47.9
GSM8K 77.9 76.7 76.5 77.9 0.0 0.0
LLaMA-3B HellaSwag 70.4 70.7 70.0 70.4 26.2 26.5
TriviaQA 33.8 31.6 27.6 33.8 0.0 0.0
TruthfulQA 49.7 49.9 49.9 49.7 51.5 48.2
GSM8K 64.6 65.4 62.2 64.6 0.0 0.0
Qwen-7B HellaSwag 80.4 80.0 80.0 80.4 79.5 77.6
TriviaQA 32.4 33.7 35.3 32.4 27.7 21.8
TruthfulQA 64.7 63.6 63.2 64.7 63.4 56.5
GSM8K 81.9 82.1 83.6 81.9 65.9 1.1
Building on our previous findings that SGR-Edit consistently outperforms its counterparts in both single and sequential medical editing, we further examine whether the edited LLMs can preserve their general capabilities. We evaluate them on four representative benchmarks: HellaSwag (commonsense reasoning), TriviaQA (factual recall), TruthfulQA (truthfulness) and GSM8K (mathematical reasoning).
RQ5-F1: Sequential-based KE method largely preserves general LLM capabilities, whereas the traditional KE method causes severe degradation and even model collapse.
As shown in Table 2, the enhanced sequential editing AlphaEdit preserves most general capabilities even after 100 sequential edits, remaining a more reliable solution for sustained medical knowledge updating while preserving the original LLM capabilities. In contrast, MEMIT exhibits drastic degradation—its edited LLaMA models collapse completely on knowledge- and reasoning-intensive tasks (TriviaQA and GSM8K drop to 0.0). Upon inspection, these edited LLaMA models produce unreadable or nonsensical outputs (e.g., random punctuation), a failure pattern also reported in prior literature (Gu et al., 2024; Yang et al., 2024). This model collapse phenomenon is not unique to MEMIT. In Appendix D.11, we present an extended analysis comparing AlphaEdit, ROME, and MEMIT across three editing paradigms (SGR, RE, GTA). The results confirm that traditional parameter-modifying methods (ROME, MEMIT) consistently suffer from catastrophic collapse in sequential settings, whereas AlphaEdit remains robust regardless of the editing paradigm.
@@ 6 Conclusion
This study introduces MedEditBench, a rigorous evaluation framework for medical knowledge editing designed to measure the reliability and generalization of updated knowledge in medical contexts. Our comprehensive evaluation reveals that the prevailing knowledge editing paradigm editing only simple final answers, often results in superficial memorization, failing to foster the deep internalization of medical principles necessary for generalization. To address this limitation, we propose the novel KE paradigm called SGR-Edit, which targets the LLM’s evidence-based reasoning chain instead of just the final answer. Extensive experiments demonstrate that SGR-Edit significantly boosts KE performance. Furthermore, our systematic analysis of sequential editing shows that the advanced KE method can effectively mitigate the degradation of broader knowledge and general abilities. These findings validate SGR-Edit as a robust KE strategy and provide crucial practical guidance for the sustainable deployment of explainable KE in real-world applications.
@@
  • Model Diversity. We only evaluate two LLaMA-3 variants (3B, 8B) and Qwen2.5-7B in this study. Since our framework applies to any LLM by targeting its failure cases, broader validation on diverse and reasoning-enhanced models (e.g., Qwen3 (Yang et al., 2025a)) is left for future work.
@@
Acknowledgments
In this work, S. Pan was supported by an Australian Research Council (ARC) Future Fellowship (FT210100097) and DP240101547.